35g of water at 0 degree Celsius is completely converted to ice at 0 degree Celsius by extracting 2 kilo joules of heat from it. If the specific heat capacity of water is 4200 joules per kilogram per Kelvin, calculate the specific latent heat of fusion of ice


1
Expert's answer
2021-03-09T15:27:07-0500

The amount of heat extracting from the water can be written as follows:


"Q=Q_1+Q_2,"

here, "Q_1" is the amount of heat that should be extracted to cool the water from "10^{\circ}C" to "0^{\circ}C", "Q_2" is the amount of heat that should be extracted to convert water from "0^{\circ}C" to ice at "0^{\circ}C".

Then, we can write:


"Q=mc\Delta T+mL_f."

From this equation we can find the specific latent heat of fusion of ice:


"L_f=\dfrac{Q-mc\Delta T}{m},""L_f=\dfrac{2\cdot10^3\ J-0.035\ kg\cdot4200\ \dfrac{J}{kg\cdot \!^{\circ}C}\cdot10^{\circ}C}{0.035\ kg}=1.5\cdot10^4\ \dfrac{J}{kg}."